OpenStudy (muscrat123):

CAN ANYBODY HELP ME WITH A FEW QUESTIONS? I WILL FAN AND MEDAL!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

sure

OpenStudy (muscrat123):

thanku thank u

OpenStudy (anonymous):

yeah

OpenStudy (muscrat123):

are u good with math?

OpenStudy (muscrat123):

i really need to get an A

OpenStudy (anonymous):

yep

OpenStudy (muscrat123):

ok hold on. i need to start bc i didnt want to start until i knew i had someone to help

OpenStudy (anonymous):

sure

OpenStudy (muscrat123):

The two dot plots below compare the forearm lengths of two groups of schoolchildren: Based on the visual inspection of the dot plots, which group appears to have the longer average forearm length? Group X, because no child in the group has a forearm length longer than 12 inches Group X, because four children in the group have the least forearm length of 10 inches Group Y, because a child in the group has the least forearm length of 10 inches Group Y, because six children have a forearm length longer than 12 inches

OpenStudy (anonymous):

A or group X

OpenStudy (anonymous):

B sorry

OpenStudy (anonymous):

Group X, because four children in the group have the least forearm length of 10 inches is answer

OpenStudy (muscrat123):

the answer is b ?

OpenStudy (anonymous):

yes

OpenStudy (muscrat123):

What is the initial value of the function represented by this table? x y 0 3 1 5 2 7 0 2 3 5

OpenStudy (anonymous):

oh i got it, hope this help: 0 + 3 =3 1+ 4 =5 2+5 =7 3+6=9 5+7=12

OpenStudy (anonymous):

so B

OpenStudy (anonymous):

@Leong

OpenStudy (muscrat123):

why b?

OpenStudy (anonymous):

Because he answer increase by two each time

OpenStudy (anonymous):

the I mean

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

:0

OpenStudy (anonymous):

next question?

OpenStudy (anonymous):

@muscrat123 look at my answers, you see how i got the results? now look at the number i plus in :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

it increase with the results on side y! tadaaaaaaa

OpenStudy (muscrat123):

yay thx

OpenStudy (anonymous):

any more questions?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

you chose correct answer :) great

OpenStudy (muscrat123):

Two functions, P and Q, are described as follows: Function P y = 5x + 3 Function Q The slope is 2 and the y-intercept is 4. How much more is the slope of function P than the slope of function Q? 3 5 7 8

OpenStudy (anonymous):

@Leong

OpenStudy (anonymous):

@_greatmath7 I have to disagree with your answer on "The two dot plots below compare the forearm lengths of two groups of schoolchildren: Based on the visual inspection of the dot plots, which group appears to have the longer average forearm length?"

OpenStudy (muscrat123):

yea actually now i think it is d. thx for the catch! @heretohelpalways

OpenStudy (muscrat123):

@Leong @LynFran

OpenStudy (muscrat123):

@robtobey

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I have to do homework I am so so sorry I feel so bad :(

OpenStudy (muscrat123):

only 3 more questions!

OpenStudy (anonymous):

gtg I am sorry

OpenStudy (anonymous):

ok 1 more question

OpenStudy (muscrat123):

really fast

OpenStudy (anonymous):

oh and it is D my bad LOL

OpenStudy (muscrat123):

Two functions, P and Q, are described as follows: Function P y = 5x + 3 Function Q The slope is 2 and the y-intercept is 4. How much more is the slope of function P than the slope of function Q? 3 5 7 8

OpenStudy (anonymous):

@iwanttolearn1212

OpenStudy (anonymous):

@hep_people

OpenStudy (anonymous):

@Lena772

OpenStudy (muscrat123):

can u help me real fast? @_greatmath7

OpenStudy (anonymous):

I will try But I don't really know functions

OpenStudy (muscrat123):

There is a 90% chance of rain tomorrow and a 30% chance that the baseball game will be postponed. What is the probability that it will rain tomorrow and the baseball game will be postponed? 3% 6% 27% 60%

OpenStudy (anonymous):

P(A & B) = P(A) * P(B) = .90 * .30 = .27

OpenStudy (anonymous):

So C

OpenStudy (muscrat123):

positive?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I am great at percents and fractions

OpenStudy (anonymous):

go to desmond graphing calc and type in both questions and you will get answer

OpenStudy (anonymous):

@muscrat123

OpenStudy (muscrat123):

for the last question i posted? which question?

OpenStudy (anonymous):

the last question you posted

OpenStudy (muscrat123):

how do i use it?

OpenStudy (anonymous):

here let me put link

OpenStudy (anonymous):

or math papa

OpenStudy (anonymous):

use math papa

OpenStudy (anonymous):

you type in y=-3x-2 in algebra calc

OpenStudy (anonymous):

how do you put links

OpenStudy (muscrat123):

this is my last question: Two functions, P and Q, are described as follows: Function P y = 5x + 3 Function Q The slope is 2 and the y-intercept is 4. How much more is the slope of function P than the slope of function Q? 3 5 7 8

OpenStudy (anonymous):

someone help @Loser66 @Leong

OpenStudy (anonymous):

WILL GIVE MEDAL AND BE FAN PLEASE HELP

OpenStudy (muscrat123):

@Zale101

OpenStudy (anonymous):

@iwanttolearn1212

OpenStudy (muscrat123):

@Owlcoffee

OpenStudy (muscrat123):

Two functions, P and Q, are described as follows: Function P y = 5x + 3 Function Q The slope is 2 and the y-intercept is 4. How much more is the slope of function P than the slope of function Q? 3 5 7 8

OpenStudy (muscrat123):

will fan

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (muscrat123):

@kropot72

OpenStudy (anonymous):

@jammy987

OpenStudy (muscrat123):

@Owlcoffee r u gonna reply?

OpenStudy (muscrat123):

@shabbaranks

OpenStudy (muscrat123):

@Katherine2016

OpenStudy (muscrat123):

Two functions, P and Q, are described as follows: Function P y = 5x + 3 Function Q The slope is 2 and the y-intercept is 4. How much more is the slope of function P than the slope of function Q? 3 5 7 8

OpenStudy (anonymous):

@Kaelyn78

OpenStudy (owlcoffee):

Function "P" represents a line, you can determine what geometric body it forms by looking at the variable with the highest exponent, and since both x and y have an exponent of "1" we can conclude that funciton "P" represents a line. A line is all the set of points that form a colineal vector with \(\vec u\) which is called the "director vector" of the line, and the line is composed by all the points that form a colineal vector with \(\vec u \) on the axis defined by the direction of the vector and any given point. That+s pretty much all there is to "lines", but what is most important is that lines can be respresented with equation that follow a certain structure, there are a total of 8 forms to express the equation of a line, but you only need to worry about two of them; The general form of the line : \(ax+by+c=0\) and the "point-slope form": \(y=mx+c\) You can see that the line represented by the function "P" follows the structure of the second structure: \(y=mx+c\) where "m" is the slope and "c" is the y-interception. This means, that only observing the equation of the function you can pull out the slope and the y-interception of the function "P". \[y=5x+3\] We can see that it follows the structure \(y=mx+c\) so therefore, we can see that the slope of function "P" is 5 and the y-intercept being "3".

OpenStudy (muscrat123):

i already figured it out but thanks

OpenStudy (anonymous):

sorry I'd really like to help but I'm not practiced enough to be confident that I can teach anyone

OpenStudy (anonymous):

for this question: 5-2